A car starts from rest and moves on a surface in which the co efficient of friction between the road and tyres increases linearly with distance (x) .The car moves with maximum possible acceleration .How will the kinetic energy of the car depend on (x)?

Grade:12th pass

1 Answers

Arun
6 years ago

 

Dear Preeti

\mu = k*x , a = \mu*g = k*x*g

v = u +at = 0 + k*x*g*t

K.E. = E = (1/2) m* v² = (1/2)* m* k²x²g²t²

Proportional to x²

Regards

Arun (askIITians forum expert)

If found useful, please approve it.

In case of any query, feel free to ask.